OpenStudy (dmnts):

sin x = cos 15 + cos 75

OpenStudy (fibonaccichick666):

are you allowed a calculator?

OpenStudy (dmnts):

I need formulas to go from cos to sin

OpenStudy (shubhamsrg):

cos a + cos b = 2 cos ((a+b)/2) cos ((a-b)/2)

OpenStudy (fibonaccichick666):

well you could just sin^-1 the whole thing, that's the easy way

OpenStudy (anonymous):

cos(theta)= sin (90-theta) so cos 15=sin 75

OpenStudy (fibonaccichick666):

or use @shubhamsrg idea or @gorvs

OpenStudy (anonymous):

@DMNTS follow me

OpenStudy (fibonaccichick666):

I prefer @gorvs

OpenStudy (anonymous):

now sinx= sin75+cos75 divide and multiply right side by\[\sqrt{2}\]

OpenStudy (anonymous):

sinx=\[\sin x=\sqrt{2}*(1/\sqrt{2}*\sin 75+1/\sqrt{2}*\cos75)\]

OpenStudy (anonymous):

1/\[1/\sqrt{2}=\sin 45 = \cos 45\]

OpenStudy (anonymous):

\[sinx= \sin75*\cos45+\sin45*\cos75\]

OpenStudy (anonymous):

sry root 2 is also there in multiplication

OpenStudy (dmnts):

Thnx, now I know how to go from here

OpenStudy (anonymous):

\[\sin x=\sqrt{2}* \sin(75+45)\]

OpenStudy (anonymous):

@DMNTS okkk buddy